# Subexponential lower bounds for $f$-ergodic Markov processes

### v1

@@ 6. Proofs of the examples in Section 3
It is thus sufficient to show that $\mathbb{P}_{x}(\{\underline{\theta}_{k}\leq S_{(\ell)}\leq t\wedge\overline{\theta}_{k+1}\}\cap A_{t})=0$ for every $k\in{\mathbb{N}}$. By (49), on $A_{t}$, the negative jumps of $X$ can only come from the Lévy process $L^{(M)}$. The modulus of negative jumps of $L^{(M)}$ on the event $A_{t}$ is bounded above by $1/4$, implying $X_{\underline{\theta}_{k}}\geq\exp(\ell+1)-1/4$. Clearly, for $u\in[\underline{\theta}_{k},\overline{\theta}_{k+1})$, we have $X_{u}\leq\exp(\ell+2)$. Hence, by (48), on the event $A_{t}$ we obtain
$\inf_{\underline{\theta}_{k}\leq s\leq t\wedge\overline{\theta}_{k+1}}X_{s}$ $=X_{\underline{\theta}_{k}}+\inf\left\{-\mu\int_{\underline{\theta}_{k}}^{s}X_{u}{\mathrm{d}}u+L_{s}^{(M)}-L_{\underline{\theta}_{k}}^{(M)}+L_{s}^{(+)}-L_{\underline{\theta}_{k}}^{(+)}:\underline{\theta}_{k}\leq s\leq t\wedge\overline{\theta}_{k+1}\right\}$
@@
where the second inequality follows from $t<1/(2\mu\exp(\ell+2))$. Hence, on the event $A_{t}$, the process $X$ cannot go below $\exp(\ell)$ during the time interval $[\underline{\theta}_{k},t\wedge\overline{\theta}_{k+1}]$ for any $k\in{\mathbb{N}}$, implying $\mathbb{P}_{x}(\{\underline{\theta}_{k}\leq S_{(\ell)}\leq t\wedge\overline{\theta}_{k+1}\}\cap A_{t})=0$ for all $k\in{\mathbb{N}}$. By (51) the claim follows.
@@
To conclude the proof of the proposition, pick $r\in(\ell+1,\infty)$ and $x\in\{\ell+1\leq V<r\}$. Note that by the claim above we have $\{S_{(\ell)}>T^{(r)}\}\supset\{S_{(\ell)}>t\}\cap\{T^{(r)}<t\}\cap A_{t}=\{T^{(r)}<t\}\cap A_{t}$ $\mathbb{P}_{x}$-a.s. for any $t\in(0,1/(2\mu\exp(\ell+2)))$. Since the positive jumps of $L$, greater than one, can only come from $L^{(+)}$ in (48), we have
$\{T^{(r)}<t\}\cap A_{t}$$\supset\left\{\sup_{0\leq s\leq t}(L_{s}-L_{s-})\geq V^{-1}(r)\right\}\cap A_{t}=\left\{\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq V^{-1}(r)\right\}\cap A_{t}$
(recall $V^{-1}(r)=\exp(r)>\exp(\ell_{0})>1$). Since the Lévy process $L^{(+)}$ is independent of $A_{t}$, we get
$\mathbb{P}_{x}(T^{(r)}<S_{(\ell)})$ $\geq\mathbb{P}_{x}(\{T^{(r)}<t\}\cap A_{t})\geq\mathbb{P}\left(\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq\exp(r)\right)\mathbb{P}(A_{t})$
$=\left(1-\exp(-t\nu(\exp(r),\infty))\right)\mathbb{P}(A_{t})>\nu(\exp(r),\infty)t\mathbb{P}(A_{t})/2$
(recall $t\nu(\exp(r),\infty))<t\nu(\exp(\ell_{0}),\infty))<\nu(\exp(\ell_{0}),\infty))/(2\mu\exp(\ell_{0}+2))<1$) for the last inequality). Proposition 6.5 follows from (50) and 3.2 (lower bound on the tail of $\nu$). ∎
@@
Recall that the solution $X$ of (17) is a strong Markov process, all the skeletons are irreducible with respect to the Lebesgue measure and compact sets are petite [34, Lem. 1.1, Prop. 1.2].
@@ Acknowledgements
MB and AM are supported by EPSRC grant EP/V009478/1. AM is also supported by EPSRC grant EP/W006227/1.

### v2

@@ 6. Proofs of the examples in Section 3
It is thus sufficient to show that $\mathbb{P}_{x}(\{\underline{\theta}_{k}\leq S_{(\ell)}\leq t\wedge\overline{\theta}_{k+1}\}\cap A_{t})=0$ for every $k\in{\mathbb{N}}$. By (49), on $A_{t}$, the negative jumps of $X$ can only come from the Lévy process $L^{(M)}$. The modulus of the negative jumps of $\int_{0}^{\cdot}\sigma(X_{s-}){\mathrm{d}}L^{(M)}_{s}$ on the event $A_{t}$ is by definition (49) bounded above by $1/4$, implying $X_{\underline{\theta}_{k}}\geq\exp(\ell+1)-1/4$. Clearly, for $u\in[\underline{\theta}_{k},\overline{\theta}_{k+1})$, we have $X_{u}\leq\exp(\ell+2)$. Hence, by (48), on the event $A_{t}$ we obtain
$\inf_{\underline{\theta}_{k}\leq s\leq t\wedge\overline{\theta}_{k+1}}X_{s}$ $=X_{\underline{\theta}_{k}}+\inf_{s\in[\underline{\theta}_{k},t\wedge\overline{\theta}_{k+1}]}\left\{-\mu\int_{\underline{\theta}_{k}}^{s}X_{u}{\mathrm{d}}u+\int_{\underline{\theta}_{k}}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)}+\int_{\underline{\theta}_{k}}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(+)}\right\}$
@@
where the second inequality follows from $t<1/(2\mu\exp(\ell+2))$. Hence, on the event $A_{t}$, the process $X$ cannot go below $\exp(\ell)$ during the time interval $[\underline{\theta}_{k},t\wedge\overline{\theta}_{k+1}]$ for any $k\in{\mathbb{N}}$, implying $\mathbb{P}_{x}(\{\underline{\theta}_{k}\leq S_{(\ell)}\leq t\wedge\overline{\theta}_{k+1}\}\cap A_{t})=0$ for all $k\in{\mathbb{N}}$. By (50), the claim follows.
@@
To conclude the proof of the proposition, pick $r\in(\ell+1,\infty)$ and $x\in\{\ell+1\leq V<r\}$. Note that by Claim 1 above we have $\{S_{(\ell)}>T^{(r)}\}\supset\{S_{(\ell)}>t\}\cap\{T^{(r)}<t\}\cap A_{t}=\{T^{(r)}<t\}\cap A_{t}$ $\mathbb{P}_{x}$-a.s. for any $t\in(0,1/(2\mu\exp(\ell+2)))$. Since the positive jumps of $L$ greater than one can only come from $L^{(+)}$ in (48), we have
$\{T^{(r)}<t\}\cap A_{t}\supset\left\{\sup_{0\leq s\leq t}(L_{s}-L_{s-})\geq\frac{V^{-1}(r)}{\sigma_{-}}\right\}\cap A_{t}=\left\{\sigma_{-}\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq V^{-1}(r)\right\}\cap A_{t}$
(recall $V^{-1}(r)=\exp(r)>\exp(\ell_{0})>1$). For some $t,c\in(0,\infty)$, Claim 2 below yields
$\mathbb{P}_{x}(T^{(r)}<S_{(\ell)})$ $\geq\mathbb{P}_{x}(\{T^{(r)}<t\}\cap A_{t})\geq\mathbb{P}_{x}\left(\{\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq\exp(r)/\sigma_{-}\}\cap A_{t}\right)$
$\geq c\left(1-\exp(-t\nu(\exp(r)/\sigma_{-},\infty))\right)>c\nu(\exp(r)/\sigma_{-},\infty)t/2,$
where the constant $c\in(0,\infty)$ is such that the third inequality holds uniformly in $r\in(\ell+1,\infty)$ and $x\in\{V\geq\ell+1\}$. Recall
$t\nu(\exp(r)/\sigma_{-},\infty))<t\nu(\exp(\ell_{0})/\sigma_{-},\infty))<\nu(\exp(\ell_{0})/\sigma_{-},\infty))/(2\mu\exp(\ell_{0}+2))<1$
for the last inequality. Proposition 6.5 now follows from the lower bound in 3.2.
Claim 2. There exists $c\in(0,\infty)$ such that
$\mathbb{P}_{x}(\{\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq\exp(r)/\sigma_{-}\}\cap A_{t})>c(1-\exp(-t\nu(\exp(r)/\sigma_{-},\infty)))$
for all $x\in\{V\geq\ell+1\}$, $r\in(\ell+1,\infty)$ and all sufficiently small $t\in(0,\infty)$.
Proof of Claim 2. The processes $L^{(-)}$ and $L^{(+)}$ are independent with $\mathbb{P}_{x}(L_{t}^{(-)}=0)=e^{-t\nu((-\infty,-1])}$ and $\mathbb{P}_{x}(\sup_{0\leq s\leq t}(L_{s}^{(+)}-L_{s-}^{(+)})\geq\exp(r)/\sigma_{-})=1-\exp(-t\nu(\exp(r)/\sigma_{-},\infty))$ for all $t\in{\mathbb{R}}_{+}$. For any $t>0$, let $\mathcal{B}_{t}$ be $\sigma$-algebra generated by $(L^{(+)}_{s},L^{(-)}_{s})_{s\in[0,t]}$. We have to show that there exists a constant $\widetilde{c}\in(0,\infty)$ such that $\mathbb{P}_{x}(\sup_{0\leq s\leq t}\left|\int_{0}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)}\right|\leq 1/8\big{|}\mathcal{B}_{t})>\widetilde{c}$, for all small $t\in(0,\infty)$ and $x\in\{V\geq\ell+1\}$. By Markov’s inequality, it is sufficient to identify a constant $C\in(0,\infty)$, such that $\operatorname{\mathbb{E}}_{x}[\sup_{0\leq s\leq t}|\int_{0}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)}|^{2}\big{|}\mathcal{B}_{t}]\leq Ct$ holds for all $t\in(0,\infty)$ and $x\in\{V\geq\ell+1\}$.
For any $y\in{\mathbb{R}}$, let the process $(Y_{t}^{y})_{t\in{\mathbb{R}}_{+}}$ be the unique strong solution[28, Thm V.6] of the SDE
(51) $Y_{t}^{y}=y-\mu\int_{0}^{t}Y_{s}^{y}{\mathrm{d}}s+\int_{0}^{t}\sigma(Y_{s-}^{y}){\mathrm{d}}L_{s}^{(M)},\quad\text{for all $t\in{\mathbb{R}}_{+}$.}$
For any pair of deterministic sequences $S=(s_{k})_{k\in{\mathbb{N}}}$ and $J=(j_{k})_{k\in{\mathbb{N}}}$, such that $0<s_{k}\uparrow\infty$ and $j_{k}\in{\mathbb{R}}\setminus\{0\}$, define the process $Z$ recursively as follows:
$Z^{S,V}_{s}\coloneqq Y_{s}^{x}{{\mathbbm{1}}\mkern-1.5mu}{\{s\in[0,s_{1})\}}+\sum_{k\in{\mathbb{N}}}{{\mathbbm{1}}\mkern-1.5mu}{\{s\in[s_{k},s_{k+1})\}}Y^{Z^{S,V}_{s_{k}}+j_{k}}_{s-s_{k}}.$
Note that the conditional law $(\int_{0}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)})_{s\in[0,t]}$, given the $\sigma$-algebra $\mathcal{B}_{t}$, equals the law of $(\int_{0}^{s}\sigma(Z^{S,V}_{u-}){\mathrm{d}}L_{u}^{(M)})_{s\in[0,t]}$ for the sequences $S$ and $V$ equal to the jump times and sizes of the components of the Lévy process $(L^{(+)}_{s},L^{(-)}_{s})_{s\in[0,t]}$. This is because the Lévy-driven SDE in (51) has a unique strong solution. Since $\sigma$ is bounded, [28, Thm V.66] yields a constant $C\in(0,\infty)$ such that for all $x\in{\mathbb{R}}$ and any pair of deterministic sequences $S$ and $J$ satisfying the above conditions, we have $\operatorname{\mathbb{E}}[\sup_{0\leq s\leq t}|\int_{0}^{s}\sigma(Z^{S,V}_{u-}){\mathrm{d}}L_{u}^{(M)}|^{2}]\leq Ct$ for all $t\in{\mathbb{R}}_{+}$. This implies that
$\operatorname{\mathbb{E}}_{x}\left[\sup_{0\leq s\leq t}\left|\int_{0}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)}\right|^{2}\Big{|}\mathcal{B}_{t}\right]\leq Ct\quad\text{for all $t\in{\mathbb{R}}_{+}$ and $x\in{\mathbb{R}}$,}$
which concludes the proof of Claim 2. ∎
@@
The process $X$ is a Feller process by [19, Thm 1.1], since we can represent it as a solution of the SDE ${\mathrm{d}}X_{t}=x+\int_{0}^{t}\widetilde{\sigma}(X_{s-}){\mathrm{d}}\widetilde{L}_{s}$ for the Lévy process $\widetilde{L}_{t}=(t,L_{t})$ and a covariance matrix $\widetilde{\sigma}(x)=(-\mu x,\sigma(x))^{\intercal}$ (note that the condition on the Lévy measure of [19, Thm 1.1] holds since $\sigma$ is bounded). Moreover, irreducibility of $X$ follows from the irreducibility of the Lévy driver $L$ in (16). Furthermore, Harris recurrence and ergodicity follow by applying [9, Thm 3.2] with a Lyapunov function $g_{2}$ and using an argument analogous to that in [12, Sec. 3.1].
@@
Recall that the solution $X$ of (17) is a strong Markov process, all the skeletons are irreducible with respect to the Lebesgue measure and compact sets are petite [38, Lem. 1.1, Prop. 1.2].
@@ Acknowledgements
We thank Martin Hairer for useful comments that improved the presentation of the paper and for bringing references [15, 37] to our attention. MB and AM are supported by EPSRC grant EP/V009478/1. AM is also supported by EPSRC grant EP/W006227/1.